OpenStudy (anonymous):

If two lines intersect and create two angles adjacent to each other which measure 10x + 8 and 5x + 7, what is the value of x?

terenzreignz (terenzreignz):

 When adjacent angles are formed by intersecting lines, their sum is always 2pi, right? :)

terenzreignz (terenzreignz):

nope, it's not 2pi, it's just pi sorry my bad :)

OpenStudy (campbell_st):

well if the angles are adjacent then they are also supplementary...add to 180 the the sum of the 2 angles is 10x + 8 + 5x + 7 = 180 15x + 15 = 180 subtract 15 from both sides 15x = 165 I'll let you solve for x

OpenStudy (anonymous):

thank you @campbell_st

OpenStudy (campbell_st):

hope it helps

OpenStudy (campbell_st):

if it is 2 lines intersecting then adjacent angles will be supplementary... just look at your diagram terenzreignz

terenzreignz (terenzreignz):

Yeah, I know that, never mind... :)

OpenStudy (anonymous):

campbell was right

OpenStudy (jiteshmeghwal9):

As we know that the sum of two adjacent angles is 180 degrees. so, the equation will be:- (10x+8)+(5x+7)=180. Now solve it & get ur answer.